Following are three sample equations. Plot them on a graph in which Q is on the vertical
axis and P is on the horizontal axis. Then transform these equations so P is expressed in
terms of Q and plot these transformed equations on a graph in which P is on the vertical
axis and Q is on the horizontal axis.
a. Q = 250 - 10P
b. Q = 1,300 - 140P
c. Q = 45 - 0.5P
1
Expert's answer
2020-11-03T11:27:42-0500

a. Q = 250 - 10P or P = 25 - 0.1Q

b. Q = 1,300 - 140P or P = 65/7 - Q/140,

c. Q = 45 - 0.5P or P = 90 - 2Q.

These equations represent the demand curves which are downward-sloping lines either P is on vertical or on horizontal axis.